OpenStudy (anonymous):

Convert 288 hrs to yr

OpenStudy (anonymous):

.0328549

OpenStudy (anonymous):

can you please shoe me the units you had to use

OpenStudy (espex):

\[288 hrs * \frac{1day}{24hr} * \frac{1yr}{365day}\]

OpenStudy (anonymous):

To be precise, a year in fact has 365.25 days

OpenStudy (anonymous):

thanks @eSpeX can you help me with a few more?

OpenStudy (espex):

Post them in the main forum and I will help if I can.

OpenStudy (anonymous):

alrighty thanks

OpenStudy (anonymous):

wait @eSpeX the final answer will be 0.03287671233

OpenStudy (espex):

Yes, unless you would like to use the more technically correct formula \[288 hours * \frac{1 day}{24 hours} * \frac{1 year}{365.25 days}\] then your answer will be 0.03285

OpenStudy (anonymous):

Who knows what the teacher wants? Just a suggestion...

OpenStudy (espex):

It's good to be technically correct, you are less likely to get an answer wrong. :)